OpenStudy (anonymous):

Determine the equation(s) of the asymptotes: g(x)=3x^2-12 ----------- x+4

OpenStudy (anonymous):

(3x^2-12)/(x+4)=3(x^2-4)/(x+4)=3(x-4)(x+4)/(x+4)=3(x-4)=3x-4